OpenStudy (anonymous):

Which polynomial function best represents the four ordered pairs given in the example? Reminder: The ordered pairs are (1, 6); (2, 11); (4, 39); and (7, -54).

OpenStudy (anonymous):

A. x3 + 8x2 + 10x - 1 = f(x) B. x3 - 4x2 + 10x - 1 = f(x) C. -2x3 + 17x2 - 32x + 23 = f(x) D. -2x3 - 17x2 - 23x + 32 = f(x)